# Multi-GPU Quantum Circuit Simulation and the Impact of Network Performance

### v1

@@ Abstract
Quantum computing , State vector , Simulation , GPU , CUDA-Q , MPI , Interconnectorganization=NVIDIA,city=Santa Clara, state=CA, country=USA
organization=Davidson School of Chemical Engineering, Purdue University,city=West Lafayette, state=IN, country=USA
@@
organization=NVIDIA,city=Santa Clara, state=CA, country=USA
organization=Davidson School of Chemical Engineering, Purdue University,city=West Lafayette, state=IN, country=USA
@@ 1 Introduction
Quantum computing has seen significant advances over the last decade, with qubit counts sufficient to run algorithms of interest, the realization of the first logical qubit for fault-tolerant computing [bluvstein2024logical], and the first claims of quantum advantage [google2025observation]. The development of new quantum algorithms capable of exploiting the advantages of quantum computing is as important as advances in the hardware itself. While quantum computers are now more capable, classical simulation of these quantum algorithms remains critical to advancing the field. Experiments on quantum computers are time-consuming, with many challenges resulting from the need for calibration and noise mitigation. Fault-tolerant quantum computers might require orders of magnitude more physical qubits to achieve reliable error correction, and we expect simulation to continue to play a critical role well into the fault-tolerant era. Of course, simulation is still used in pre-silicon assessment and post-silicon debugging for the development of new classical computers. Those who have been involved in the supercomputing race will be very familiar with the need to validate bleeding-edge hardware and software based on expected behavior, a challenge amplified in quantum computing due to the wide variety of novel engineering designs.
@@
Rigorous performance evaluation is critical as quantum systems transition from research prototypes to practical computational tools. Benchmarking efforts span multiple levels of the quantum computing stack, from component-level hardware characterization [Proctor2025-cd, PracticalCharacter, cross2019validating, qiskit_measuring_quantum_volume, baldwin2022re, pelofske2022quantum] and system-level performance evaluation [proctor2022measuring, blume2020volumetric, proctor2022establishing, wack_clops_2021, wack2110quality] to compiler and toolchain measurement [kharkov2022arline, nation2025benchmarking] and algorithmic performance analysis [Supermarq, QASMbench, lubinski2023_10061574, Quark, Boixo_2018, bernal2025benchmarking]. While each provides valuable insights within its focus area, application-oriented benchmarks are essential for evaluating end-to-end system performance on representative workloads. In this work, we utilize the QED-C Application-Oriented Performance Benchmarks for Quantum Computing [lubinski2023_10061574, lubinski2023optimization, lubinski2024quantumalgorithmexplorationusing, chatterjee2024comprehensivecrossmodelframeworkbenchmarking, niu2025practicalframeworkassessingperformance, QEDC-App-Benchmarks], an evolving suite of quantum algorithms and applications designed to assess performance on real-world problems. To enable scalable benchmarking on HPC platforms, we introduce support for MPI [MPI] into the QED-C framework for distributed multi-GPU evaluation. This implementation-agnostic approach supports performance assessment across quantum programming frameworks, providing insights relevant to both near-term simulation and the development of fault-tolerant quantum systems
@@
GPU-acceleration can provide dramatic speedups for classical simulation and has become standard practice [vallero2025state, faj2023quantum] with popular simulation frameworks such as Qiskit [javadi2024quantum], PennyLane [bergholm2018pennylane], Cirq [cirq], and CUDA-Q [kim2023cuda] supporting GPUs. At maximum GPU memory capacity, we typically measure speedups of 2-3 orders of magnitude when comparing state-of-the-art CPUs and GPUs on a single socket, due to better-optimized algorithms and higher peak hardware throughput. While techniques for gate fusion [smelyanskiy2016qhipster, zhang2021hyquas, faj2023quantum, bayraktar2023cuquantum] can result in near-peak utilization of main memory bandwidth (BW) and floating-point throughput when using a single GPU, network communications can become the bottleneck when distributing a simulation across multiple GPUs. This is relevant as state-vector simulation, the most generally applicable simulation approach, scales exponentially in memory and computational requirements with the qubit count.
@@
For 32-bit floating-point precision, the memory requirement for the state vector in gigabytes (GiB) is given by for qubits. With this encoding, current GPUs are limited to simulating around 34 qubits (128 GiB state vector) when constrained by a single GPU’s memory. While host memory can be used to supplement storage, the need for significant HPC resources can become necessary, where multiple GPUs across multiple nodes are required either to attain the aggregate memory requirements for a given qubit count or to achieve reasonable time-to-solution for the simulation of many circuits (for example, with iterative algorithms, noisy simulation, AI-guided applications, etc.). In these cases, the performance of data movement becomes a concern as data is transferred between distributed memory domains across different GPUs or potentially between the host memory and the GPU.
@@
The interconnect technologies for network communications between GPUs have undergone significant advancements since the introduction of GPU acceleration to HPC. While the first systems required a path involving the host memory, support for remote direct memory access (RDMA), first marketed by NVIDIA as “GPUDirect” was introduced to allow direct memory access between multiple GPUs on the same node or from a network interface controller (NIC) for internode communications [gpudirect]. To overcome the limitations of the PCI Express bus initially used for intranode multi-GPU communications, high-bandwidth GPU-specialized interconnects were first introduced under the marketing name NVLink [nvlink] (NVL) (Infinity Fabric for AMD Systems). With NVLink 4, systems combining NVIDIA Grace CPUs with GPUs feature a specialized coherent interconnect between the CPU and GPU, providing the same peak bidirectional bandwidth. This is called NVLink-C2C. While these specialized interconnects have been limited to communications within a node, NVIDIA has recently released Grace Blackwell NVL72 [nvl72] as the first generally available system to expand this high-BW all-to-all interconnect across multiple nodes, referred to here as multi-node NVL (MNNVL). In the case of NVL72, up to 72 GPUs across multiple nodes can be in the same NVL domain.
@@
| MI350X Infinity Fabric[mi300] | 153.6 GB/s | X | |||
| Slingshot 11 [perlmutter] | 25 GB/s | X | |||
@@ 2.1 Terminology
Classical simulation of quantum algorithms commonly relies on the state-vector representation, in which the full quantum state of an -qubit system is stored as a complex vector of dimension . Each quantum gate is applied through a tensor product of an appropriate unitary operator. The exponential scaling of the state-vector size imposes significant computational and memory demands.
@@ 2.2 Benchmarks
In this work, we use the quantum phase estimation (QPE) [shukla2025practicalquantumphaseestimation] benchmark as it represents a fundamental algorithmic component in quantum computing applications and is amenable to weak-scaling studies due to the simple scaling behavior of gate count with the number of qubits. For a more sophisticated benchmark, we also evaluate strong scaling using a 33-qubit Transverse-field Ising model [TFIM] Hamiltonian from HamLib [Sawaya2024hamliblibraryof] with 2D periodic boundary conditions on a triagonal lattice.
@@
The QPE circuit consists of two main components: (i) a register of qubits on which controlled powers of unitary gate are applied (e.g., ), and (ii) an inverse Quantum Fourier Transform (QFT) that extracts the phase information into the computational basis. The QED-C [lubinski2024quantum] implementation samples phases that are exactly representable with qubits (i.e., ), enabling clean fidelity comparisons. Increasing increases both circuit width and depth while preserving algorithmic regularity. This makes QPE a valuable reference for evaluating circuit-generation time, transpilation overhead, execution time, and fidelity as a function of system size. Within the QED-C framework, average execution times, circuit depths, and fidelity between the measured and ideal phase distributions are recorded as key performance metrics.
@@
HamLib [Sawaya2024hamliblibraryof] provides a standardized set of Hamiltonians, including condensed matter physics models like the Heisenberg model, the Fermi-Hubbard model, and the Transverse-field Ising model (TFIM), chemistry models of molecules such as and , and combinatorial optimization models. The library is designed for reproducible benchmarking of quantum algorithms and different backends. The size of the Hamiltonians ranges from a few qubits to large-scale instances with hundreds or even thousands of qubits.
In this work, we study the time evolution of the TFIM [TFIM] implemented using Suzuki-Trotter [SUZUKI1990319, Dragoi_Trotter] decomposition, producing deep circuits with nontrivial interaction structure that closely resemble realistic quantum-simulation workloads. The TFIM is used extensively in benchmarking quantum-simulation workloads. The Hamiltonian for a system of particle spins is given by
@@ 2.3 Quantum Simulation Software
We evaluate performance using NVIDIA CUDA-Q [cudaq], a framework for heterogeneous quantum-classical computing. CUDA-Q put forward a programming model, in C++ and Python, comprising types, concepts, syntax, and semantics, which facilitates the integration of quantum processing units (QPUs) and high-performance simulation backends into conventional computing workflows.
In CUDA-Q, QPU code is defined as a quantum kernel, which is annotated for a custom compilation pipeline, based on MLIR, to produce an executable payload for the quantum coprocessor/simulator backend. There are various hardware QPUs and simulators available in CUDA-Q. Specifically, both state vector and tensor network-based simulators, based on the cuQuantum library [bayraktar2023cuquantum], are available in CUDA-Q. Each backend, either QPU or simulator, is assigned a unique target name, allowing users to interchange backends in a hardware-agnostic manner.
@@ 2.4 APIs for multi-GPU Communication
The recommended approach for exploiting multi-GPU parallelism is with high-level libraries for inter-GPU communication. These include GPU-aware MPI, libraries based on SHMEM [bariuso1994], or newer libraries that have been heavily optimized for AI workloads. For NVL72, the newest releases of Unified Communication X [shamis2015ucx] (UCX) underlying MPI, NVSHMEM, and NVIDIA Collective Communications Library (NCCL) have all added optimized support. Additionally, these libraries can be used in conjunction with each other; both NVSHMEM and NCCL can be used along with MPI in the same application. For “ninja” level optimizations, lower-level APIs can be used for finer control of inter-GPU communications. For GPUDirect and NVLink, these include the CUDA interprocess communication (IPC) API and the virtual memory management (VMM) API available through the CUDA Driver. For AMD GPUs, the direct peer GPU memory access functionality is exposed by replacing the characters “cuda” with “hip” for functions in the NVIDIA-designed IPC API. For MNNVL, the IPC API is not supported, and the VMM API should be used.
Quantum circuit simulators have been implemented using a PGAS model supporting NVSHMEM [li2021sv], and MNNVL is designed to be efficient with a global address space across GPUs. AI-optimized libraries, such as NCCL, can offer performance benefits due to both business prioritization and the potential freedom from optimization limitations imposed by other standards. Both of these APIs can be particularly beneficial where fine-grain communication can be exploited for performance without involvement of the GPU host. Nonetheless, MPI remains most prevalent in HPC and is the high-level API used for internode communication in cuQuantum and CUDA-Q.
@@
Note that full error handling for has been omitted from these examples and that even when running on only a single node, the cuMemCreate function can fail if the system does not have properly configured IMEX channels [imex]. An alternative approach to the VMM API is to allocate MPI buffers using stream-ordered memory allocators (cudaMallocAsync(), cudaFreeAsync()) with a memory pool that has been created with the CU_MEM_HANDLE_TYPE_FABRIC property (see cudaMemPoolCreate()). In addition to CUDA-runtime support, this also carries the typical advantages for stream-ordered allocation, including low overhead where it is difficult to avoid allocation within a loop.
@@ 2.5 Multi-node Benchmarking and MPI
MPI support in cuQuantum facilitates both decreasing time to solution for a given system size or increasing the system size beyond what will fit in the memory of a single GPU or node. Both are achieved by distributing the state vector into smaller sub-statevectors across multiple GPUs to exploit increased memory capacity, BW, and compute throughput [bayraktar2023cuquantum]. As this distribution requires intensive communication that can occur between any pairs of GPUs (see [jones2023distributed for details on communication patterns), we demonstrate here that network communications can quickly become a bottleneck as scaling occurs.
cuQuantum supports supplementing GPU memory with host memory; however, it is expected that this feature will come with a performance impact. For example, when simulating with two additional qubits using host memory, one can expect a fourfold increase in simulation time compared to using only GPU memory. This is simply because the simulation is performed on of the number of GPUs. In practice, the performance hit can be significantly worse, as all the data in host memory must be moved to the GPU for computation at a bandwidth that may be limited by host memory itself or the bus to the GPU. While this feature is certainly valuable for fully exploiting the limits of a given system or job size, here we limited the scope to focus on GPU scalability for high performance.
@@
The QED-C Application-Oriented Benchmark Suite is an open-source framework designed to facilitate community contributions and extensibility. The modular architecture supports multiple quantum programming APIs, including Qiskit and CUDA-Q, enabling straightforward integration of new capabilities such as the MPI functionality described here. Here, we describe the implementation details of the MPI extensions to the Python-based benchmarks, with complete source code available in the QED-C Application-Oriented Benchmarks repository [QEDC-App-Benchmarks].
@@
A final important implementation note is the use of synchronization for benchmarking. That is, we enforce a barrier before starting timing so that statistics are not influenced by implicit synchronization over variable startup costs or conditional execution. MPI support has been added to a subset of the QED-C benchmarks, and we are currently expanding.

### v2

@@ Abstract
Quantum computing , State vector , Simulation , GPU , CUDA-Q , MPI , Interconnect[a]organization=NVIDIA,address=2788 San Tomas Expressway, city=Santa Clara, state=CA, zip=95051, country=USA
[b]organization=Davidson School of Chemical Engineering, Purdue University,address=480 Stadium Mall Dr, city=West Lafayette, state=IN, zip=47907, country=USA
@@ 1 Introduction
Quantum computing has seen significant advances over the last decade, with qubit counts sufficient to run algorithms of interest, the realization of the first logical qubit for fault-tolerant computing [9], and the first claims of quantum advantage [21]. The development of new quantum algorithms capable of exploiting the advantages of quantum computing is as important as advances in the hardware itself. While quantum computers are now more capable, classical simulation of these quantum algorithms remains critical to advancing the field. Experiments on quantum computers are time-consuming, with many challenges resulting from the need for calibration and noise mitigation. Fault-tolerant quantum computers might require orders of magnitude more physical qubits to achieve reliable error correction, and we expect simulation to continue to play a critical role well into the fault-tolerant era. Of course, simulation is still used in pre-silicon assessment and post-silicon debugging for the development of new classical computers. Those who have been involved in the supercomputing race will be very familiar with the need to validate bleeding-edge hardware and software based on expected behavior, a challenge amplified in quantum computing due to the wide variety of novel engineering designs.
@@
Rigorous performance evaluation is critical as quantum systems transition from research prototypes to practical computational tools. Benchmarking efforts span multiple levels of the quantum computing stack, from component-level hardware characterization [48, 23, 12, 58, 3, 45] and system-level performance evaluation [46, 8, 47, 64, 63] to compiler and toolchain measurement [28, 37] and algorithmic performance analysis [61, 31, 35, 18, 10, 7]. While each provides valuable insights within its focus area, application-oriented benchmarks are essential for evaluating end-to-end system performance on representative workloads. In this work, we utilize the QED-C Application-Oriented Performance Benchmarks for Quantum Computing [35, 32, 34, 11, 39, 49], an evolving suite of quantum algorithms and applications designed to assess performance on real-world problems. To enable scalable benchmarking on HPC platforms, we introduce support for MPI [36] into the QED-C framework for distributed multi-GPU evaluation. This implementation-agnostic approach supports performance assessment across quantum programming frameworks, providing insights relevant to both near-term simulation and the development of fault-tolerant quantum systems.
@@
GPU-acceleration can provide dramatic speedups for classical simulation and has become standard practice [62, 17] with popular simulation frameworks such as Qiskit [24], PennyLane [6], Cirq [14], and CUDA-Q [29] supporting GPUs. At maximum GPU memory capacity, we typically measure speedups of 2-3 orders of magnitude when comparing state-of-the-art CPUs and GPUs on a single socket, due to better-optimized algorithms and higher peak hardware throughput. While techniques for gate fusion [55, 67, 17, 5] can result in near-peak utilization of main memory bandwidth (BW) and floating-point throughput when using a single GPU, network communications can become the bottleneck when distributing a simulation across multiple GPUs. This is relevant as state-vector simulation, the most generally applicable simulation approach, scales exponentially in memory and computational requirements with the qubit count.
@@
For 32-bit floating-point precision, the memory requirement for the state vector in gigabytes (GiB) is given by for qubits. With this encoding, current GPUs are limited to simulating around 34 qubits (128 GiB state vector) when constrained by a single GPU’s memory. While host memory can be used to supplement storage, the need for significant HPC resources can become necessary, where multiple GPUs across multiple nodes are required either to attain the aggregate memory requirements for a given qubit count or to achieve reasonable time-to-solution for the simulation of many circuits (for example, with iterative algorithms, noisy simulation, AI-guided applications, etc.). Significant work has gone into the research and implementation of efficient multi-GPU algorithms for distributed quantum circuit simulation [40, 13, 68, 55, 22, 26, 30, 67, 5, 25, 66, 65, 19, 59, 50]. In these cases, the performance of data movement becomes a concern as data is transferred between distributed memory domains across different GPUs or potentially between the host memory and the GPU.
@@
The interconnect technologies for network communications between GPUs have undergone significant advancements since the introduction of GPU acceleration to HPC. While the first systems required a path involving the host memory, support for remote direct memory access (RDMA), first marketed by NVIDIA as “GPUDirect” was introduced to allow direct memory access between multiple GPUs on the same node or from a network interface controller (NIC) for internode communications [41]. To overcome the limitations of the PCI Express bus initially used for intranode multi-GPU communications, high-bandwidth GPU-specialized interconnects were first introduced under the marketing name NVLink [42] (NVL) (Infinity Fabric for AMD Systems). With NVLink 4, systems combining NVIDIA Grace CPUs with GPUs feature a specialized coherent interconnect between the CPU and GPU, providing the same peak bidirectional bandwidth. This is called NVLink-C2C. While these specialized interconnects have been limited to communications within a node, NVIDIA has recently released Grace Blackwell NVL72 [43] as the first generally available system to expand this high-BW all-to-all interconnect across multiple nodes, referred to here as multi-node NVL (MNNVL). In the case of NVL72, up to 72 GPUs across multiple nodes can be in the same NVL domain.
@@
| MI350X Infinity Fabric [2] | 153.6 GB/s | X | |||
| Slingshot 11 [38] | 25 GB/s | X | |||
@@ 2.1 Terminology
Classical simulation of quantum algorithms commonly relies on the state-vector representation, in which the full quantum state of an -qubit system is stored as a complex vector of dimension . Each quantum gate is applied through a tensor product of an appropriate unitary operator. In general, gate application typically dominates the simulation run time and due to the space complexity for the state vector, the time complexity in the general case is with gate count . The exact number of floating point operations and the associated memory access patterns depend on the type of gate, the number of qubits involved, and the index of the gate within the circuit. Therefore, the simulation times at fixed qubit and gate counts can be expected to vary depending on the specifics of the circuit. When distributing a state vector across memory domains, the number of elements exchanged between domains scales as for the simulation in the general case; some gate applications can be achieved without exchanging any elements with remote domains. As there is sufficient local memory to store the state vector for qubits for memory domains, distributed algorithms can partition the circuit such that gate applications involving this subset do not require inter-GPU communications. In addition to the qubit index and number of memory domains, it is also the case that the exact number of elements exchanged will depend on the type of gate.
For a single-qubit gate, for example, gate applications involve multiplying state vector elements with a 2x2 matrix defined by the gate. Given the small dimension, it is expected that this operation would be memory-BW bound at sufficient qubit counts on most system architectures. Optimization with gate fusion can be employed to group multiple gate applications into a single operation. This does not necessarily reduce floating point operations and can actually result in some increase. However, fusion can increase linear algebra dimensions, improve memory reuse at the cache and register levels, and shift the gate applications from BW-bound to a balanced arithmetic intensity. In this case, the performance goal does not involve maximizing the fusion level, but rather tuning it for the target architecture such that the simulation is not bottlenecked by memory or compute throughput. Thus, the optimal fusion level can depend on factors such as the GPU architecture and the floating-point precision used for simulation.
@@ 2.2 Benchmarks
In this work, we use the quantum phase estimation (QPE) [54] benchmark as it represents a fundamental algorithmic component in quantum computing applications and is amenable to weak-scaling studies due to the simple scaling behavior of gate count with the number of qubits. For a more sophisticated benchmark, we also evaluate strong scaling using a 33-qubit Transverse-field Ising model [16] Hamiltonian from HamLib [52] with 2D periodic boundary conditions on a triagonal lattice. As a final benchmark, we simulate random circuit samples (RCS) to better cover the general case by including irregular circuits with random connectivity.
@@
The QPE circuit consists of two main components: (i) a register of qubits on which controlled powers of unitary gate are applied (e.g., ), and (ii) an inverse Quantum Fourier Transform (QFT) that extracts the phase information into the computational basis. The QED-C [33] implementation samples phases that are exactly representable with qubits (i.e., ), enabling clean fidelity comparisons. Increasing increases both circuit width and depth while preserving algorithmic regularity. This makes QPE a valuable reference for evaluating circuit-generation time, transpilation overhead, execution time, and fidelity as a function of system size. Within the QED-C framework, average execution times, circuit depths, and fidelity between the measured and ideal phase distributions are recorded as key performance metrics.
@@
HamLib [52] provides a standardized set of Hamiltonians, including condensed matter physics models like the Heisenberg model, the Fermi-Hubbard model, and the Transverse-field Ising model (TFIM), chemistry models of molecules such as and , and combinatorial optimization models. The library is designed for reproducible benchmarking of quantum algorithms and different backends. The size of the Hamiltonians ranges from a few qubits to large-scale instances with hundreds or even thousands of qubits.
In this work, we study the time evolution of the TFIM [16] implemented using Suzuki-Trotter [57, 15] decomposition, producing deep circuits with nontrivial interaction structure that closely resemble realistic quantum-simulation workloads. The TFIM is used extensively in benchmarking quantum-simulation workloads. The Hamiltonian for a system of particle spins is given by
@@
Random Circuits
For benchmarking of random circuit samples (RCS), we use a simple Python script outside of the QED-C benchmark suite that starts with an empty circuit of qubits and loops to add gates on randomly chosen qubits. Here, the probability of a two-qubit controlled ”Z” gate is 0.25, and otherwise single-qubit gates are chosen from ”H”, ”RX”, ”RY”, and ”RZ” as defined in the CUDA-Q documentation [60]. Each random circuit is simulated twice, with only the second simulation timed. This allows us to exclude just-in-time compilation of the CUDA-Q kernels, as they are cached, and compilation is not benchmarked as part of this work. A fixed random seed is used to generate consistent circuits for benchmarking in different configurations.
The unfused gate count for the HamLib benchmark was 2987. For QPE, the gate count ranged from 625 at 33 qubits to 856 at 39 qubits. For RCS, the gate counts were for qubits. Default CUDA-Q settings for gate fusion were used in this work, which generally give optimal performance at higher qubit counts and resulted in the best performance for these benchmarks. For FP32, the default fusion levels are 4, 5, and 5 for Ampere, Hopper, and Blackwell, respectively. There is currently no option to determine the number of gates after fusion (and therefore, this would require instrumenting the code). To normalize ideal parallel scaling performance, we provide approximate ideal numbers that assume a uniform reduction in gate counts from fusion across different qubit/GPU counts.
@@ 2.3 Quantum Simulation Software
We evaluate performance using NVIDIA CUDA-Q [60], a framework for heterogeneous quantum-classical computing. CUDA-Q put forward a programming model, in C++ and Python, comprising types, concepts, syntax, and semantics, which facilitates the integration of quantum processing units (QPUs) and high-performance simulation backends into conventional computing workflows.
In CUDA-Q, QPU code is defined as a quantum kernel, which is annotated for a custom compilation pipeline, based on MLIR, to produce an executable payload for the quantum coprocessor/simulator backend. There are various hardware QPUs and simulators available in CUDA-Q. Specifically, both state vector and tensor network-based simulators, based on the cuQuantum library [5], are available in CUDA-Q. Each backend, either QPU or simulator, is assigned a unique target name, allowing users to interchange backends in a hardware-agnostic manner.
@@ 2.4 APIs for multi-GPU Communication
The recommended approach for exploiting multi-GPU parallelism is with high-level libraries for inter-GPU communication. These include GPU-aware MPI, libraries based on SHMEM [4], or newer libraries that have been heavily optimized for AI workloads. For NVL72, the newest releases of Unified Communication X [53] (UCX) underlying MPI, NVSHMEM, and NVIDIA Collective Communications Library (NCCL) have all added optimized support. Additionally, these libraries can be used in conjunction with each other; both NVSHMEM and NCCL can be used along with MPI in the same application. For “ninja” level optimizations, lower-level APIs can be used for finer control of inter-GPU communications. For GPUDirect and NVLink, these include the CUDA interprocess communication (IPC) API and the virtual memory management (VMM) API available through the CUDA Driver. For AMD GPUs, the direct peer GPU memory access functionality is exposed by replacing the characters “cuda” with “hip” for functions in the NVIDIA-designed IPC API. For MNNVL, the IPC API is not supported, and the VMM API should be used.
Quantum circuit simulators have been implemented using a PGAS model supporting NVSHMEM [30], and MNNVL is designed to be efficient with a global address space across GPUs. AI-optimized libraries, such as NCCL, can offer performance benefits due to both business prioritization and the potential freedom from optimization limitations imposed by other standards. Both of these APIs can be particularly beneficial where fine-grain communication can be exploited for performance without involvement of the GPU host. Nonetheless, MPI remains most prevalent in HPC and is the high-level API used for internode communication in cuQuantum and CUDA-Q.
@@
Note that full error handling for has been omitted from these examples and that even when running on only a single node, the cuMemCreate function can fail if the system does not have properly configured IMEX channels [44]. An alternative approach to the VMM API is to allocate MPI buffers using stream-ordered memory allocators (cudaMallocAsync(), cudaFreeAsync()) with a memory pool that has been created with the CU_MEM_HANDLE_TYPE_FABRIC property (see cudaMemPoolCreate()). In addition to CUDA-runtime support, this also carries the typical advantages for stream-ordered allocation, including low overhead where it is difficult to avoid allocation within a loop.
@@ 2.5 Multi-node Benchmarking and MPI
MPI support in cuQuantum facilitates both decreasing time to solution for a given system size or increasing the system size beyond what will fit in the memory of a single GPU or node. Both are achieved by distributing the state vector into smaller sub-statevectors across multiple GPUs to exploit increased memory capacity, BW, and compute throughput [5]. As this distribution requires intensive communication that can occur between any pairs of GPUs (see [27, 13] for details on communication patterns), we demonstrate here that network communications can quickly become a bottleneck as scaling occurs.
With parallel state-vector simulation in cuQuantum, it is required that the number of GPUs used for a single circuit be a power of 2. cuQuantum supports supplementing GPU memory with host memory; however, it is expected that this feature will come with a performance impact. For example, when simulating with two additional qubits using host memory, one can expect a fourfold increase in simulation time compared to using only GPU memory. This is simply because the simulation is performed on of the number of GPUs.111In practice, one would only supplement with host memory when there is insufficient GPU memory available. Due to the memory requirements when adding qubits and the restriction that the number of GPUs involved in parallelization must be a power of 2, at least four times as many GPUs would be required to fit the simulation entirely in GPU memory when adding 2 qubits. In practice, the performance hit can be significantly worse, as all the data in host memory must be moved to the GPU for computation at a bandwidth that may be limited by host memory itself or the bus to the GPU. While this feature is certainly valuable for fully exploiting the limits of a given system or job size, here we limited the scope to focus on GPU scalability for high performance.
@@
The QED-C Application-Oriented Benchmark Suite is an open-source framework designed to facilitate community contributions and extensibility. The modular architecture supports multiple quantum programming APIs, including Qiskit and CUDA-Q, enabling straightforward integration of new capabilities such as the MPI functionality described here. Here, we describe the implementation details of the MPI extensions to the Python-based benchmarks, with complete source code available in the QED-C Application-Oriented Benchmarks repository [49].
@@
Care must be taken with MPI parallelization to assert that random number generation produces the same result on all MPI tasks in cases where these numbers can impact the simulated circuit and parametrization. In general, it is expected that all tasks invoke CUDA-Q calls with the same parameters, and behavior otherwise is undefined and might be difficult to debug at runtime. A final important implementation note is the use of synchronization for benchmarking. That is, we enforce a barrier before starting timing so that statistics are not influenced by implicit synchronization over variable startup costs or conditional execution.
Our inclusion of stubs for parallel operations in the QED-C benchmarks, along with use of the mpi4py Python module and cuQuantum plugins for MPI support, allows for simple interoperability with CUDA-Q across systems with or without MPI or different implementations of MPI prior to ABI standardization. Additionally, the approach should ease the transition to future support for alternative communication libraries. We have added support for a subset of the QED-C benchmarks and are expanding. Support is currently limited to CUDA-Q. With expanded support, another concern is interoperability across different simulation frameworks. While we expect this to be largely transparent for frameworks using cuQuantum as a backend, multi-GPU parallelization choices for other backends could force more significant changes. However, we expect these to be minor compared with the larger effort required to consistently generate quantum circuits across frameworks with significant differences in programming, compilation, and execution models.